OpenStudy (anonymous):

hey guys I need some help and I am not sure if I post it here or in science but I will give it a try I am doing this lab and in the lab I measured how long it would take for a sound wave to get to five meters well I timed it several times and I got 190 in seconds my teacher wants to use this v=d/t but I don't know how to could someone explain it to me

OpenStudy (anonymous):

this means velocity = distance/time

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

yes sorry my computer was having problems would I divide 190/5=38 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sorry, thats upside down. 5/190

OpenStudy (anonymous):

okay so 5/190= 0.0263157895

OpenStudy (anonymous):

yeah and remember your units. if its a velocity its probably m/s

OpenStudy (anonymous):

okay thank you your amazing do you think you could help me with another thing I am tying to find but I cant

OpenStudy (anonymous):

im not sure if you did that right though because the speed of sound at sea level is 340.29 m / s

OpenStudy (anonymous):

yeah, i can help with another question

OpenStudy (anonymous):

i got it at http://phet.colorado.edu/en/simulation/sound okay i think there is a different way to calculate this but i don't know how

OpenStudy (anonymous):

do you know or is there just one way

OpenStudy (anonymous):

i got .0182 seconds when i did it

OpenStudy (anonymous):

never mind i found another way frequency=velocity/wavelength did you play around with this lab or did you find the calculations for what i gave to you

OpenStudy (anonymous):

played around with it for a bit

OpenStudy (anonymous):

oh see my teacher played with it to he said some people would get a different answer he got 0.130 when he did it

OpenStudy (anonymous):

well either way its definitely smaller than 190

OpenStudy (anonymous):

okay i will do it again

OpenStudy (anonymous):

oh wait did you stop it when it got to five and not when it got to the end of the stick

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh now i feel stupid you were right sorry about that

OpenStudy (anonymous):

okay see now i got 0.172

OpenStudy (anonymous):

great, that makes alot more sense. now plug that in instead

OpenStudy (anonymous):

5/0.172=29.0697674m/s

OpenStudy (anonymous):

is that right